# Worse than Random? An Embarrassingly Simple Probing Evaluation of Large Multimodal Models in Medical VQA

### v1

@@ A.12 Abdominal X-ray
Random Choice acc 25 25 50 36.55 37.46
acc w.o. adv. pair 50 50 / 37.54 37.51
LLaVA-v1-7B acc 53.87 53.01 50 2.15 0.56
acc w.o. adv. pair 100 99.56 / 95.25 99.45
LLaVA-v1.6-7B acc 5.17 56.46 46 6.46 1.12
acc w.o. adv. pair 97.84 78.44 / 80.17 91.52
LLaVA-Med acc 7.75 42.24 60 14.65 4.51
acc w.o. adv. pair 98.27 91.37 / 23.7 89.83
MiniGPT-v2 acc 4.74 38.79 50 18.53 5.64
acc w.o. adv. pair 99.13 86.63 / 56.03 67.23
CheXagent acc 77.15 23.7 70 12.93 2.25
acc w.o. adv. pair 90.51 46.55 / 38.79 37.85
GPT-4V acc 84.84 50.21 60 31.16 23.16
acc w.o. adv. pair 84.84 71.86 / 42.85 37.85
Gemini Pro acc 97.14 63.36 85 27.15 19.2
acc w.o. adv. pair 99.13 96.55 / 73.27 67.23
@@
Appendix B ProbMed
Table 19: Number of questions across each question type for each image. Ground-truth questions were created based on available metadata, with "yes" answers. For each ground-truth question, we also created a corresponding adversarial question by selecting random adversarial entities and assigning "no" answers. For an image showing a normal organ without abnormality, since there is no ground-truth information on the existence of the condition and position, we only construct hallucinated questions for the condition/finding question type. For an image showing abnormality, the number of question pairs per category equals the number of existing conditions or positions.

### v5

@@ A.12 Abdominal X-ray
Random Choice Acc. with adv. pairs 25 25 50 36.55 37.46
LLaVA-v1 Acc. with adv. pairs 53.87 53.01 50 2.15 0.56
Avg. acc. 76.93 76.50 / 49.14 50.00
LLaVA-v1.6 Acc. with adv. pairs 5.17 56.46 46 6.46 1.12
Avg. acc. 52.15 75.64 / 47.63 48.16
MiniGPT-v2 Acc. with adv. pairs 4.74 38.79 50 18.53 5.64
Avg. acc. 52.37 67.24 / 53.65 50.23
LLaVA-Med-v1 Acc. with adv. pairs 7.75 42.24 60 14.65 4.51
Avg. acc. 53.23 68.96 / 47.47 50.87
LLaVA-Med-v1.5 Acc. with adv. pairs 52.58 50.86 50.00 6.46 14.68
Avg. acc. 76.07 73.49 / 52.02 54.75
BiomedGPT Acc. with adv. pairs 35.77 1.29 50.00 10.34 4.51
Avg. acc. 65.30 37.50 / 52.94 46.79
Med-Flamingo Acc. with adv. pairs 28.01 34.48 50 14.65 4.51
Avg. acc. 64.00 66.37 / 52.52 46.25
CheXagent Acc. with adv. pairs 77.15 23.70 70 12.93 2.25
Avg. acc. 88.57 52.80 / 51.30 49.64
GPT-4o Acc. with adv. pairs 98.26 61.47 70 27.27 21.46
Avg. acc. 99.13 79.22 / 61.83 59.81
GPT-4V Acc. with adv. pairs 84.84 50.21 60 31.16 23.16
Avg. acc. 92.42 71.42 / 59.63 57.03
Gemini Pro Acc. with adv. pairs 97.14 63.36 85 27.15 19.2
Avg. acc. 98.70 80.81 / 59.97 58.80
@@
Appendix B Dataset Statistics
In this section, we detail the dataset statistics.
Table 17 shows the number of questions across each question type. For each image sample, ground-truth questions were created based on available metadata with "yes" answers. For each ground-truth question, we also created a corresponding adversarial question by selecting random adversarial entities and assigning "no" answers. For an image showing a normal organ without abnormality, since there is no ground-truth information on the existence of the condition and position, we only construct hallucinated questions for the condition/finding question type. For an image showing abnormality, the number of question pairs per category equals the number of existing conditions or positions.
Table 17: Number of questions across each question type for each image.
@@ Appendix B Dataset Statistics
Appendix C Robustness Study
We conducted an additional robustness analysis using GPT-4o—the best-performing model from our evaluations—on the subset of 100 images and 1,090 VQA pairs used in our expert validation. The following table presents results on randomly sampled test samples from the expert-validated subset (20%, 40%, 60%, 80%, and 100%) and compares accuracy across all samples versus manually validated ("valid") samples:
Table 18: Valid percentage and accuracy under all samples vs. valid samples on different test sample splits. $N_{i}$ denotes the number of images in the split, and $N$ denotes the number of corresponding question-answer pairs.
$N_{i}$, $N$ Valid QA (%) Acc. all samples (%) Acc. valid samples (%)
20, 218 100.00 52.63 52.63
40, 421 96.19 54.36 55.19
60, 596 96.81 55.32 55.25
80, 821 95.73 54.21 54.55
100, 1090 97.79 55.63 55.59
6303, 57132 / 55.60 /
From these results, we observe minimal differences between the accuracy of all samples versus the validated subset. This indicates our evaluation framework is highly robust to the approximately 6% error rate identified during pre-processing. Furthermore, the accuracy of GPT-4o on the full ProbMed dataset (55.60%) closely aligns with the performance on the subset (55.63%), confirming the generalizability of our robustness findings.
Appendix D Impact of Chain-of-Thought Prompts and Visual Descriptions on Model Performance
In this section, we illustrate model performance under different categories under the three prompting settings: Vanilla, Chain-of-Thought, and Chain-of-Thought with GPT-4o Visual Understanding.
Figure 6: Accuracy of the LLaVA-v1 model across five diagnostic categories under three settings: vanilla (blue), chain-of-thought (CoT, red), and CoT with GPT-4o Visual Understanding (green).
Figure 7: Accuracy of the LLaVA-v1.6 model across five diagnostic categories under three settings: vanilla (blue), chain-of-thought (CoT, red), and CoT with GPT-4o Visual Understanding (green).
Figure 8: Accuracy of the LLaVA-Med-v1 model across five diagnostic categories under three settings: vanilla (blue), chain-of-thought (CoT, red), and CoT with GPT-4o Visual Understanding (green).
Figure 9: Accuracy of the LLaVA-Med-v1.5 model across five diagnostic categories under three settings: vanilla (blue), chain-of-thought (CoT, red), and CoT with GPT-4o Visual Understanding (green).
Figure 10: Accuracy of the Med-Flamingo model across five diagnostic categories under three settings: vanilla (blue), chain-of-thought (CoT, red), and CoT with GPT-4o Visual Understanding (green).
Figure 11: Accuracy of the GPT-4o model across five diagnostic categories under three settings: vanilla (blue), chain-of-thought (CoT, red), and CoT with GPT-4o Visual Understanding (green).
Appendix E Prompt Details
Below is the few-shot model prompt used for GPT-4 to extract medical conditions and their locations from image captions:
1You are a helpful assistant and you are given a caption describing a medical image. Extract medical conditions and diseases, along with their locations, if specified. Do not include any information that cannot be directly inferred from the image, for example, patient status or patient history. Outputs should be in the format: "<condition/disease1> : <location1>, <condition/disease2> : <location2>...". The term "<location>" should include at least one positional descriptor and should be explicit in the original caption along with the condition/disease. Otherwise, it should be replaced with "None".
2
3For example, consider the caption: "Fig. 1. MRI abdomen and pelvis showing the cervical mass." The output should be "<cervical mass> : None". For the caption: "Chest radiograph shows enlargement of the hilar mass with spread into the left lower lobe." The output should be "<enlargement of the hilar mass> : <left lower lobe>". Similarly, for the caption: "Abdominal CT scan reveals an enhancing rounded pseudo-aneurysm in the cystic artery, alongside high-density material within the gallbladder’s lumen and near the gastrohepatic ligament." The correct output is "<enhancing rounded pseudo-aneurysm> : <cystic artery>, <high-density material> : <lumen of the gallbladder and region of the gastrohepatic ligament>".
4
5Make sure that the response contains only the information in the original caption without adding extra details.
Appendix F Response Distribution Visualization within each Category
Below, we illustrate the distribution of model responses within each category, valid "yes" responses are in red, and valid "no" responses are in blue.
Figure 12: Distribution plot of "yes and "no" ground-truth answers and model responses within the Modality category.
Figure 13: Distribution plot of "yes and "no" ground-truth answers and model responses within the Organ category.
Figure 14: Distribution plot of "yes and "no" ground-truth answers and model responses within the Abnormality category.
Figure 15: Distribution plot of "yes and "no" ground-truth answers and model responses within the Condition/Finding category.
Figure 16: Distribution plot of "yes and "no" ground-truth answers and model responses within the Position category.